OpenStudy (anonymous):

I seriously need help! Tutor please? ABC has vertices A (0, 0), B (2, 2), and C (5, -1). Find the coordinates of L, the midpoint of AC, and M, the midpoint of BC. Verify that LM AB and LM= 1/2 AB.

OpenStudy (anonymous):

And please explain

OpenStudy (anonymous):

@Hero @HelloKitty17 @jakyfraze @texaschic101

OpenStudy (aravindg):

First can you try to draw a figure for better understanding?

OpenStudy (anonymous):

you must use the midpoint formula

OpenStudy (anonymous):

Yeah 

OpenStudy (anonymous):

isnt that y=mx+b?

OpenStudy (anonymous):

no

OpenStudy (aravindg):

See if (a,b) , (c,d) are 2 of the vertices. Then the midpoint is given by ( (a+c)/2, (b+d)/2 )

OpenStudy (anonymous):

m= y2-y1/x2-x1? But there is no d coordinate

OpenStudy (anonymous):

http://cs.selu.edu/~rbyrd/math/midpoint/mid_ex1.gif

OpenStudy (anonymous):

this is the formula you must plug in your own numbers

OpenStudy (anonymous):

I think I got it. L (5/2, -½), M (7/2, ½). AB slope = 1 and LB slope = 1 also. So LM= ½ AB

OpenStudy (anonymous):

Is that right?